# https://gaoxiangjun.github.io/mani_gs/

### v1

@@ Abstract
Neural 3D representations such as Neural Radiance Fields (NeRF), excel at producing photo-realistic rendering results but lack the flexibility for manipulation and editing which is crucial for content creation. Previous works have attempted to address this issue by deforming a NeRF in canonical space or manipulating the radiance field based on an explicit mesh. However, manipulating NeRF is not highly controllable and requires a long training and inference time. With the emergence of 3D Gaussian Splatting (3DGS), extremely high-fidelity novel view synthesis can be achieved using an explicit point-based 3D representation with much faster training and rendering speed. However, there is still a lack of effective means to manipulate 3DGS freely while maintaining rendering quality. In this work, we aim to tackle the challenge of achieving manipulable photo-realistic rendering. We propose to utilize a triangular mesh to manipulate 3DGS directly with self-adaptation. This approach reduces the need to design various algorithms for different types of Gaussian manipulation. By utilizing a triangle shape-aware Gaussian binding and adapting method, we can achieve 3DGS manipulation and preserve high-fidelity rendering after manipulation. Our approach is capable of handling large deformations, local manipulations, and soft body simulations while keeping high-quality rendering. Furthermore, we demonstrate that our method is also effective with inaccurate meshes extracted from 3DGS. Experiments conducted demonstrate the effectiveness of our method and its superiority over baseline approaches.
Keywords:
Gaussian Splatting Manipulation Rendering
@@ 1 Introduction
Figure 1: Our proposed approach allows for 3DGS manipulation, including large deformation, local manipulation, and even physical simulation (such as soft body simulation), while maintaining high-quality rendering. Please zoom in for more details.
@@
Manipulating and editing 3D content is essential for content creation and has various applications in movies, gaming, and virtual/augmented reality. 3D model editing enables users to create and modify models flexibly, thereby enhancing production efficiency. The traditional pipeline for modeling and editing a 3D asset with photo-realistic rendering involves a process with geometry modeling, texturing, UV mapping, lighting, and rendering, which is a tedious and time-consuming flow requiring lots of manual work.
@@
Over the past few years, the neural radiance field (NeRF)[21] has been widely studied due to its high capability and simple reconstruction process in 3D representation. However, the implicit representation poses challenges for editing. To address this, some methods are proposed to edit this implicit neural radiance field[21]. NeRF-Editing[43] is the first to utilize the triangular mesh to help edit the implicit radiance field. They train a canonical NeuS [31] and extract the triangular mesh from NeuS[31]. A tetrahedra grid is then constructed to contain the object mesh. To render the deformed object by editing the triangular mesh, a volume rendering is conducted in the deformed space. The sampling points in deformed space are mapped to canonical space based on the constructed tetrahedra grid, where the points in deformed space have the same tetrahedron barycentric coordinate in the same tetrahedron with their corresponding points in canonical space. Moreover, [10, 18] demonstrate similar ideas by employing tetrahedra to deform sampling points and achieve editable nerf-based scenes. Instead, Neu-Mesh[38] and SERF[47] define the neural radiance field by associating each mesh vertex with radiance and geometry features. Then they can conduct volume rendering like Point-NeRF[36] in deformed space without backward mapping. However, these editing methods based on implicit neural radiance fields still suffer from inconvenient manipulation, suboptimal rendering results, and long training and rendering times.
@@
Meanwhile, 3DGS[14] has gained significant attention in differential rendering due to its high-fidelity and fast rendering proficiency. However, despite being an explicit 3D representation, it still lacks an effective method for manipulating 3DGS while maintaining high-quality rendering. SuGaR[7], the work most closely related to our objective, develops a novel algorithm to extract a triangular mesh from 3DGS. Although their main goal is not to facilitate editable photorealistic rendering, they bind the 3DGS to the extracted mesh, enabling model animation as demonstrated in their demo.
@@
This work proposes a method that enables 3DGS manipulation, achieving high-quality and photo-realistic rendering. Our key insight is to manipulate 3DGS using a triangular mesh as the proxy, which allows for direct transfer of mesh manipulation to 3DGS with 3DGS self-adaptation. With our methods, we can achieve large deformation, local manipulation, and soft body simulaton with high-quality results as shown in Figure 1, which also avoid the need to design various algorithms for different types of manipulation.
@@
To achieve controllable 3DGS manipulation through the mesh, an intuitive approach is to bind the GS to lay perfectly on the triangle and enforce the GS to be thin enough. After mesh manipulation, the GS will automatically adapt its rotation and position with the attached triangle, as employed in SuGaR[7]. However, SuGaR heavily relies on the accuracy of mesh geometry, inheriting the defects of mesh rendering. Specifically, for inaccurate parts, SuGaR cannot inpaint the missing parts or remove the redundant parts in the final rendering.
@@
Adding an offset to the position of attached Gaussians during the reconstruction may seem like a reasonable solution to compensate for mesh inaccuracy. However, this fixed offset cannot be generalized well to the deformed space after novel manipulation. Our proposed solution is to define a local coordinate system for each triangle, which we refer to as local triangle space. We then bind Gaussians to each triangle and optimize the Gaussian attributes, including rotation, position, and scaling, in the attached local triangle space.
@@
GaMeS[6] and Mesh-GS [28] are two concurrent works that also employ triangular meshes for Gaussian Splatting manipulation. In particular, GaMeS[6] constrains the Gaussians on the surface exactly; Mesh-GS [28] permits an offset along the normal direction without adapting the Gaussian scale when the triangle shape changes. In contrast, our model allows for Gaussian move in the triangle local space which means we can achieve high-quality rendering without the need for accurate mesh. Gaussian scaling also adapts in response to changes in triangle shape when large deformations are applied to the mesh.
In summary, the contributions of our paper are listed as follows:
@@
  • We introduce a triangle shape aware Gaussian binding strategy with self-adaption, which has a high tolerance for mesh accuracy and supports various 3DGS manipulations.
  • We evaluate our method and achieve state-of-the-art results, demonstrating various 3DGS manipulations, including large deformation, local manipulation, and soft body simulation.
@@ 2.1 NeRF Editing
Recently, neural radiance field [21] (NeRF) has garnered significant attention due to its high-quality and photo-realistic rendering results for novel view synthesis. NeRF represents the scene as a continuous function that maps a spatial location and viewing direction to a volume density and color, which is parameterized by a multilayer perceptron (MLP). Owing to the implicit representation that encodes the scene within the network parameters, editing and deforming the geometry of the NeRF scene explicitly like mesh can be challenging. To enable user editing of NeRF, [19] introduce editing conditional radiance fields trained on a shape category. However, it only supports basic editing operations, such as removing/adding object parts or shape transfer. CLIP-NeRF [29] achieves NeRF editing with text or images by leveraging CLIP model [23] but still can not edit the geometry locally. Some other work[34, 32, 1, 44] edit the NeRF in texture level which is not the focus of this paper.
To edit and deform NeRF locally, [10, 43, 18] construct a tetrahedra grid based on the underlying 3D shape. After explicitly deforming the tetrahedra into the posed space for editing, the sampled 3D points are mapped from the posed space to the canonical space through the unaltered tetrahedron, which means the canonical position can be calculated from the shared barycentric coordinate for both deformed and canonical tetrahedron. The density and radiance in the posed space can be calculated for the mapped sampling points in canonical space. On the other hand, [30, 47, 38] employ mesh as the guidance for deformation. NeuMesh [38] presents a novel representation to encode neural implicit field on a mesh-based scaffold for geometry and texture editing. [30] achieves the manipulation of both the geometry and color of neural implicit fields through differentiable colored meshes. Furthermore, there are several methods [25, 20, 45, 2, 46] that particularly focus on editing NeRF for avatar. In this work, we introduce an editing method based on 3DGS for general objects.
@@ 2.2 Mesh-based NeRF Rendering
NeRFs have shown impressive rendering results, however, rendering one pixel using NeRF representation necessitates a volumetric rendering algorithm that involves inferring MLP hundreds of times to estimate their radiance and density. This process is significantly slower than traditional mesh rendering. To accelerate NeRF rendering, several methods [3, 24, 41, 3, 40, 27] have been proposed to combine NeRF representation with mesh reconstruction. By converting this implicit representation to an explicit mesh, these methods may also facilitate applications like editing. In particular, MobileNeRF [3] represents NeRF as a collection of polygons with deep feature textures, which can be rendered using the classic polygon rasterization pipeline, generating a feature vector for each pixel and passing it to an MLP to decode the color. MobileNeRF is capable of achieving rendering even on standard mobile devices. Additionally, NeRFMeshing [24] distills the reconstructed NeRF into a signed surface approximation network to extract 3D mesh and shows the simulation results for editing. BakedSDF [41] proposes a neural surface-volume representation to extract the mesh and supports editing like material decomposition, appearance editing, and physics simulation. However, the editing results of these methods heavily rely on the accuracy of the extracted mesh. Artifacts present in the mesh will directly influence the editing results. In contrast, our method demonstrates robustness to the reconstructed mesh and can still yield promising results even with the inaccurate mesh.
@@ 2.3 Gaussian Splatting Editing, Simulation and Animation
3D Gaussian Splatting [14] presents an innovative 3D Gaussian scene representation, accompanied by a differentiable renderer that attains real-time rendering of radiance fields while maintaining high quality. Initially, 3D Gaussian Splatting focuses solely on static scenes, which has been extended to model dynamic scenes [33, 9, 39, 17, 48], human avatars [49, 22, 42, 8, 37, 15], Gaussian Splatting simulation and animation [7, 35, 12, 4]. Specifically, SuGaR [7] proposes a method to extract meshes from 3D Gaussian Splatting with additional regularization, in which they bind 3DGS on extracted mesh and animated with Gaussian Splatting rendering. GSP [4] incorporates physically-based fluid dynamics in 3DGS and PhysGaussian [35] introduces a unified simulation-rendering pipeline that generates physics-based dynamics with photorealistic renderings. VR-GS [12] achieves interactive physics-based editing in Virtual Reality. In this work, we also introduce a Gaussian Splatting manipulation method that binds Gaussian Splatting to the mesh, achieving state-of-the-art results.
@@ 3 Method
Recently, due to its exceptional high-fidelity rendering capabilities and fast rendering speed, Gaussian Splatting [14] has emerged as a popular 3D representation in the differential rendering field. However, despite being an explicit 3D representation, it still lacks a methodology for manipulating this 3D representation for editing while maintaining high-quality rendering after the manipulation. In this work, giving multi-view RGB images of an object as input, we introduce a method for object manipulation that can achieve photorealistic editable rendering by employing Gaussian Splatting.
The pipeline of our method is illustrated in Figure 2 and consists of three main stages. First, we extract a mesh from 3D Gaussian Splatting (3DGS) or a neural surface field for subsequent 3D Gaussian binding (Sec. 3.2). Next, we devise a novel Mesh-Gaussian binding method dedicated to manipulating Gaussian Splatting while maintaining photo-realistic rendering quality (Sec. 3.3). Finally, we describe the types of Gaussian manipulation we support and how to render the manipulated results (Sec. 3.4).
@@ 3.2 Mesh Extraction
Our method is capable of achieving high-quality reconstruction and editing using guided meshes obtained from various methods. In this section, we investigate different mesh extraction and reconstruction techniques with different mesh accuracy and processing time to guide the 3D Gaussians in our approach.
Screened Poisson Reconstruction. 3D Gaussian Splatting could be considered a type of point cloud, making it intuitive to extract the mesh using the Poisson-reconstruction algorithm. However, the 3D Gaussians do not have normal vectors for reconstruction. Inspired by recent 3DGS inverse rendering methods [5, 16], we allocate an additional gaussian attribute, normal $\bm{n}$, for 3D Gaussians:
$\{\mathcal{D},\mathcal{N}\}=\sum_{i\in{N}}T_{i}\alpha_{i}\{d_{i},\bm{n}_{i}\},$ (4)
where $d_{i}$ denotes the depth and $\bm{n}_{i}$ denotes the normal of each gaussian point. To optimize the normal, we then encourage the rendered normal $\mathcal{N}$ to be well-aligned with the pseudo-normal map $\tilde{\mathcal{N}}$, computed from the rendered depth $\mathcal{D}$ based the local planarity assumption. After training 3DGS with normal attributes, we can extract the mesh using the Screened poisson surface reconstruction[13] algorithm.
Marching Cube for Gaussian Splatting. In DreamGaussian[26], the method attempts to summarize the alpha values of neighboring Gaussian points as the composite density value of marching cube sampling points. The Gaussian points within a pre-defined local voxel are considered neighbors. However, We found that this method often ignores the thin and small structures. To address this issue, we employ ball query Nearest Neighbor searching methods to search only the nearest Gaussian points of grid sampling points. The sampling points that have a neighbor within a pre-defined distance are assigned with a density value of $1$, otherwise with a density value of 00. The geometry is then extracted using the Marching Cube. This approach allows for the extraction of thin and small structures, even though they may be slightly larger than the actual surface. With our Mesh-Gaussian binding strategy, we can achieve high-fidelity rendering with the slightly inaccurate mesh and support smooth Gaussian manipulation.
Neural Implicit Surfaces. In this work, we also try to extract high-quality surfaces from the implicit representation utilizing the method proposed in NeuS [31]. NeuS suggests extracting a surface as the zero-level set of a signed distance function (SDF), which has consistently served as a robust and widely used benchmark in the realm of neural surface reconstruction. Nevertheless, the surface derived from NeuS may encompass an excessive number of triangle surfaces, which may have around 1.5 million. A large number of triangles can negatively affect both training and inference speeds, therefore, we employ a mesh cleaning process to eliminate the noisy floaters and utilize mesh decimation techniques to reduce the count of triangles to approximately 300K. We circumvent the need for such a high number of triangles while maintaining high-fidelity rendering.
@@ 3.3 Binding Gaussian Splatting on Mesh
For each triangle $\bm{f}$ in the given mesh $\bm{T}$, which is composed of three vertices $(\bm{v_{1}},\bm{v_{2}},\bm{v_{3}})$, we initialize $N$ Gaussians on this triangle. To be specific, the mean position $\bm{\mu}$ of initialized Gaussians is formulated as $\bm{\mu}=(w_{1}\bm{v_{1}}+w_{2}\bm{v_{2}}+w_{3}\bm{v_{3}})$, $\bm{w}=(w_{1},w_{2},w_{3})$ is the pre-defined barycentric coordinate of each Gaussians attached on the triangle. And $\bm{w}$ satisfy $(w_{1}+w_{2}+w_{3})=1$.
Gaussians on Mesh. To achieve controllable 3DGS manipulation through the mesh, an intuitive way is to perfectly attach the GS to the triangle, as shown in the SuGaR[7]. With the rotation matrix denoted as $\bm{R}=\{\bm{r_{1}},\bm{r_{2}},\bm{r_{3}}\}$ and the scaling vector represented by $\bm{s}=(s_{1},s_{2},s_{3})$, SuGaR train 3DGS with a flat Gaussian distribution on the mesh by setting $s_{1}=\epsilon$, where $\epsilon$ is close to zero. $\bm{r_{1}}$ is defined by the normal vector $\bm{n}$ of the attached triangle. The Gaussians have only 2 learnable scaling factors $(s_{2},s_{3})$ instead of 3, and only 1 learnable 2D rotation rather than a quaternion, to keep the Gaussians flat and aligned with the mesh triangles. This type of binding strategy heavily relies on mesh accuracy, which lacks the flexibility of 3DGS to model complex object rendering. Moreover, the mesh quality obtained from SuGaR is significantly inferior compared to the ground truth and recent neural surface reconstruction results, which increases the difficulty of editing.
Gaussians on Mesh with Offset. To compensate for the inaccuracy of the extracted mesh, it would be better to add an offset $\Delta\bm{\mu}$ to the Gaussians 3D mean $\bm{\mu}$, which enables the Gaussians to move out of the attached triangle $\bm{f}$. Although it could improve the rendering quality of the reconstructed static object, it would result in noisy and unexpected rendering distortion in the manipulated object due to the mismatched localized relative position between Gaussians.
Triangle Shape Aware Gaussian Binding and Adapting. To preserve the high-fidelity rendering results after manipulation, the key lies in maintaining the local rigidity and preserving the relative location between Gaussians, both for 3D means and rotations. Our key insight is to define a local coordinate system in each triangle space.
@@
where $\bm{v_{1}}$, $\bm{v_{2}}$ is the first and second vertex location repectively, $\bm{n^{t}}$ is the normal vector which can be calulated by:
@@
This Gaussian mesh binding method can preserve the relative position and rotation between Gaussians that are bonded on neighboring triangles after mesh manipulation. However, following mesh manipulation, not only does the triangle center change but also the triangle shape. With the altered triangle shape, the local Gaussian position and scaling should adjust accordingly. When the triangle enlarges, it is intuitive that the local scaling and position should expand as well:

### v2

# Mani-GS: Gaussian Splatting Manipulation with Triangular Mesh

@@ Abstract
Neural 3D representations, such as Neural Radiation Fields (NeRF), excel at producing photorealistic rendering results but lack the flexibility for manipulation and editing which is crucial for content creation. However, manipulating NeRF is not highly controllable and requires a long training and inference time. With the emergence of 3D Gaussian Splatting (3DGS), extremely high-fidelity novel view synthesis can be achieved using an explicit point-based 3D representation with much faster training and rendering speed. However, there is still a lack of effective means to manipulate 3DGS freely while maintaining rendering quality. In this work, we aim to tackle the challenge of achieving manipulable photo-realistic rendering. We propose to utilize a triangular mesh to manipulate 3DGS directly with self-adaptation. This approach reduces the need to design various algorithms for different types of 3DGS manipulation. footnotetext: Corresponding author: Xiaoyu Li and Yao Yao. By utilizing a triangle shape-aware Gaussian binding and adapting method, we can achieve 3DGS manipulation and preserve high-fidelity rendering. In addition, our method is also effective with inaccurate meshes extracted from 3DGS. Experiments demonstrate our method’s effectiveness and superiority over baseline approaches.
Figure 1: Our proposed approach allows for 3DGS manipulation, including large deformation, local manipulation, and even physical simulation (such as soft body), while maintaining high-quality rendering. (Check project page for more visual results.)
@@ 1 Introduction
Editing 3D content is essential for content creation and has various applications in movies, gaming, and virtual/augmented reality. 3D model editing enables users to create and modify models flexibly, thereby enhancing production efficiency. The traditional pipeline for modeling and editing a 3D asset with photo-realistic rendering involves a process with geometry modeling, texturing, UV mapping, lighting, and rendering, which is a tedious and time-consuming flow requiring lots of manual work.
@@
Over the past few years, the neural radiance field (NeRF) [27] has been widely studied due to its high capability and simple reconstruction process in 3D representation. However, the implicit representation poses challenges for editing. To address this, some methods are proposed to edit this implicit neural radiance field [27]. NeRF-Editing [51] is the first to utilize the triangular mesh to help edit the implicit radiance field. Volume rendering is conducted in the deformed space to render the deformed object by editing the triangular mesh. The sampling points in the deformed space are mapped to canonical space based on the constructed tetrahedra grid.  [15, 24] present similar ideas by employing tetrahedra to deform sampling points and achieve editable rendering. In addition, NeuMesh [45] and SERF [55] define the neural radiance field by associating each mesh vertex with radiance and geometry features. Then they conduct volume rendering like Point-NeRF [43] in deformed space without backward mapping. However, these editing methods based on implicit neural radiance fields still suffer from inconvenient manipulation, suboptimal rendering results, and long training and rendering times.
@@
Recently, 3DGS [19] has gained significant attention in differential rendering due to its high-fidelity and fast rendering proficiency. However, despite being an explicit 3D representation, it still lacks an effective method for manipulating 3DGS while maintaining high-quality rendering. SuGaR [10], the work most closely related to our objective, develops a novel algorithm to extract a triangular mesh from 3DGS. Although their main goal is not to facilitate editable photorealistic rendering, they bind the 3DGS to the extracted mesh, enabling model animation.
@@
This paper proposes a method that enables 3DGS manipulation, achieving high-quality and photo-realistic rendering. Our key insight is to manipulate 3DGS using a triangular mesh as the proxy, which allows for direct transfer of mesh manipulation to 3DGS with 3DGS self-adaptation. With our methods, we can achieve large deformation, local manipulation, and soft body simulaton with high-quality results as shown in Fig 1, which also avoid the need to design various algorithms for different types of manipulation.
@@
To achieve controllable 3DGS manipulation through mesh, an intuitive approach is to bind the GS to lay perfectly on the triangle and enforce the GS to be thin enough. After mesh manipulation, the GS will automatically adapt its rotation and position with the attached triangle, as employed in SuGaR [10]. However, SuGaR heavily relies on the accuracy of mesh geometry, inheriting the defects of mesh rendering. Specifically, for inaccurate parts, SuGaR cannot inpaint the missing parts or remove the redundant parts during rendering.
@@
Adding an offset to the position of attached Gaussians during the reconstruction may seem like a reasonable solution to compensate for mesh inaccuracy. However, this fixed offset cannot be generalized well to the deformed space after manipulation. Our proposed solution is to define a local coordinate system for each triangle, which we refer to as local triangle space. We then bind Gaussians to each triangle and optimize the Gaussian attributes, including rotation, position, and scaling, in the attached local triangle space.
@@
The contributions of our paper are listed as follows:
@@
  • We introduce a triangle shape aware Gaussian binding strategy with self-adaption, which has a high tolerance for mesh accuracy.
  • We evaluate our method and achieve state of the art results, also supporting various 3DGS manipulations including large deformation, local manipulation, and soft body simulation.
@@ 2.1 NeRF Editing
Recently, NeRF [27] has garnered significant attention due to its high-quality and photo-realistic rendering results for novel view synthesis. NeRF represents the scene as a continuous function that maps a spatial location and viewing direction to a volume density and color, which is parameterized by a multilayer perceptron (MLP). Owing to the implicit representation that encodes the scene within the network parameters, editing and deforming the geometry of the NeRF scene explicitly like mesh can be challenging. To enable user editing of NeRF,  [25] introduce editing conditional radiance fields trained on a shape category. However, it only supports basic editing, such as removing/adding object parts or shape transfer. CLIP-NeRF [35] achieves NeRF editing with text or images by leveraging CLIP model [29] but still can not edit the geometry locally. Some other work [41, 38, 2, 52] edit the NeRF in texture level which is not the focus of this paper.
To edit and deform NeRF locally,  [15, 51, 24] construct a tetrahedra grid based on the underlying 3D shape. After explicitly deforming the tetrahedra into the posed space for editing, the sampled 3D points are mapped from the posed space to the canonical space through the unaltered tetrahedron, which means the canonical position can be calculated from the shared barycentric coordinate for both deformed and canonical tetrahedron. The density and radiance in the posed space can be calculated for the mapped sampling points in canonical space. On the other hand,  [36, 55, 45] employ mesh as the guidance for deformation. NeuMesh [45] presents a novel representation to encode neural implicit field on a mesh-based scaffold for geometry and texture editing. [36] achieves the manipulation of both the geometry and color of neural implicit fields through differentiable colored meshes. Furthermore, there are several methods [31, 26, 53, 3, 54, 57, 40, 58, 8] that particularly focus on editing NeRF for avatar. In this work, we introduce an editing method based on 3DGS for general objects.
@@ 2.2 Mesh-based NeRF Rendering
NeRFs have shown impressive rendering results, however, rendering one pixel using NeRF necessitates a volumetric rendering algorithm that involves inferring MLP hundreds of times to estimate their radiance and density. To accelerate NeRF rendering, several methods [30, 48, 4, 47, 33, 9, 49, 12] have been proposed to combine NeRF representation with mesh reconstruction. By converting this implicit representation to an explicit mesh, these methods may also facilitate applications like editing. In particular, MobileNeRF [4] represents NeRF as a collection of polygons with deep feature textures, which can be rendered using the classic polygon rasterization pipeline. Additionally, NeRFMeshing [30] distills the reconstructed NeRF into a signed surface approximation network to extract 3D mesh and shows the simulation results for editing. However, the editing results of these methods heavily rely on the accuracy of the extracted mesh. Artifacts present in the mesh will directly influence the editing results. In contrast, our method demonstrates robustness to the reconstructed mesh and can still yield promising results even with the inaccurate mesh.
@@ 2.3 Gaussian Splatting Editing, Simulation and Animation
3D Gaussian Splatting  [19] presents an innovative 3D Gaussian scene representation, accompanied by a differentiable renderer that attains real-time rendering of radiance fields while maintaining high quality. Initially, 3D Gaussian Splatting focuses solely on static scenes, which has been extended to model dynamic scenes [39, 14, 46, 23, 56], human avatars [59, 28, 50, 13, 44, 20, 1], Gaussian Splatting simulation and animation [10, 42, 17, 5]. Specifically, SuGaR [10] proposes a method to extract meshes from 3D Gaussian Splatting with additional regularization, in which they bind 3DGS on extracted mesh and animated with Gaussian Splatting rendering. GSP [5] incorporates physically-based fluid dynamics in 3DGS and PhysGaussian [42] introduces a unified simulation-rendering pipeline that generates physics-based dynamics with photorealistic renderings. VR-GS [17] achieves interactive physics-based editing in Virtual Reality. In this work, we also introduce a 3DGS manipulation method that binds Gaussian Splatting to the mesh, achieving state-of-the-art results.
Several concurrent works have emerged in parallel with ours. In particular, GaMeS [34] constrains the 3DGS should be a flat ellipse and lay on the surface tightly, which means the position of 3DGS is inside the triangle and rotation is the same with triangle rotation; Mesh-GS  [7] permits an offset along the normal direction but the rotation and scaling are fixed after manipulation. Our model allows 3DGS to move out of the triangle and maintain the relative position and rotation after manipulation. And we can achieve high-quality rendering without the need for accurate mesh. Following SuGaR, Guedon et al. [11] proposed to represent the scene using a multi-layer mesh, while our method requires only a single-layer mesh.
@@ 3 Method
Recently, due to its exceptional high-fidelity rendering capabilities and fast rendering speed, 3DGS [19] has emerged as a popular 3D representation in the differential rendering field. However, despite being an explicit 3D representation, it still lacks a way to manipulate this 3D representation for editing while maintaining high-quality rendering after the manipulation. In this work, giving multi-view RGB images of an object as input, we introduce a method for object manipulation that can achieve photo-realistic editable rendering by employing 3DGS.
The pipeline of our method is illustrated in Figure 2 and consists of three main stages. First, we extract a mesh from 3D Gaussian Splatting (3DGS) or a neural surface field for subsequent 3D Gaussian binding (Sec. 3.2). Next, we devise a novel Mesh-Gaussian binding method dedicated to manipulating 3DGS while maintaining photo-realistic rendering quality (Sec. 3.3). Finally, we describe the types of Gaussian manipulation we support(Sec. 3.4).
@@ 3.2 Mesh Extraction
Our method can achieve high-quality editing using guided meshes obtained from various methods. In this section, we investigate different mesh extraction and reconstruction techniques with different mesh accuracy and processing time to guide the 3D Gaussians in our approach.
Marching Cube for Gaussian Splatting. In DreamGaussian [32], the method attempts to summarize the alpha values of neighboring Gaussian points as the composite density value of marching cube sampling points. The mesh extracted using this method often ignores the thin and small structures. With our Mesh-Gaussian binding strategy, we can achieve high-fidelity rendering with the inaccurate mesh and support smooth Gaussian manipulation.
Screened Poisson Reconstruction. 3D Gaussian Splatting could be considered a type of point cloud, making it intuitive to extract the mesh using the Poisson-reconstruction algorithm. However, the 3D Gaussians do not have normal vectors for reconstruction. Inspired by recent 3DGS inverse rendering methods [6, 22], we allocate an additional gaussian attribute, normal $\bm{n}$, for 3D Gaussians, which is supervised by the pseudo normal derived from depth map. After training 3DGS with normal attributes, we can extract the mesh using the Screened poisson surface reconstruction [18] algorithm.
Neural Implicit Surfaces. In this work, we also try to extract high-quality surfaces from the implicit representation utilizing the method proposed in NeuS [37]. NeuS mesh has a large number of triangles, which negatively affects both training and inference speeds. We utilize mesh decimation techniques to reduce the count of triangles to approximately 300K.
@@ 3.3 Binding Gaussian Splatting on Mesh
For each triangle $\bm{f}$ in the given mesh $\bm{T}$, which is composed of three vertices $(\bm{v_{1}},\bm{v_{2}},\bm{v_{3}})$, we initialize $N=3$ Gaussians on this triangle. To be specific, the mean position $\bm{\mu}$ of initialized Gaussians is formulated as $\bm{\mu}=(w_{1}\bm{v_{1}}+w_{2}\bm{v_{2}}+w_{3}\bm{v_{3}})$, $\bm{w}=(w_{1},w_{2},w_{3})$ is the pre-defined barycentric coordinate of each Gaussians attached on the triangle. And $\bm{w}$ satisfy $(w_{1}+w_{2}+w_{3})=1$.
Gaussians on Mesh. To achieve controllable 3DGS manipulation through the mesh, an intuitive way is to perfectly attach 3DGS to the triangle, as shown in the SuGaR [10]. With the rotation matrix denoted as $\bm{R}=\{\bm{r_{1}},\bm{r_{2}},\bm{r_{3}}\}$ and the scaling vector represented by $\bm{s}=(s_{1},s_{2},s_{3})$, SuGaR train 3DGS with a flat Gaussian distribution on the mesh by setting $s_{1}=\epsilon$, where $\epsilon$ is close to zero. $\bm{r_{1}}$ is defined by the normal vector $\bm{n}$ of the attached triangle. The Gaussians have only 2 learnable scaling factors $(s_{2},s_{3})$ instead of 3, and only 1 learnable 2D rotation rather than a quaternion, to keep the Gaussians flat and aligned with the mesh triangles. This binding strategy is heavily dependent on mesh accuracy, which limits the flexibility of 3DGS in modeling complex object rendering. When the mesh accuracy is low, it cannot compensate for missing or redundant parts of geometry. Conversely, when the mesh accuracy is high but still deviates from the ground truth, it tends to produce a blurred effect due to the view inconsistency between the inaccurate mesh and multiview images. Overall, this binding strategy inherits the shortcomings of mesh rendering.
Gaussians on Mesh with Offset. To compensate for the inaccuracy of the extracted mesh, it would be better to add an offset $\Delta\bm{\mu}$ to the Gaussians 3D mean $\bm{\mu}$, which enables the Gaussians to move out of the attached triangle $\bm{f}$. Although it can improve the rendering quality of the reconstructed static object, this offset field is not generalized. It would result in noisy and unexpected rendering distortion in the manipulated object due to the mismatched relative position between 3DGS.
Triangle Shape Aware Gaussian Binding and Adaption. To preserve the high-fidelity rendering results after manipulation, the key lies in maintaining the local rigidity and preserving the relative location between Gaussians, both for 3D means and rotations. Our key insight is to define a local coordinate system in each triangle space.
@@
where $\bm{v_{1}}$, $\bm{v_{2}}$ is the first and second vertex location repectively, $\bm{n^{t}}$ is the normal vector calulated by:
@@
Figure 3: Visual comparison between ours, NeRF-Editing [25](N.E.) and SuGaR [10] for static rendering. It illustrates our proposed method can contain a much more accurate boundary in “Ship”, and distinct results in “Drums”.
Table 1: Quantitative comparison of our methods with NeRF-Editing (“N.E.”) [51] and “SuGaR” [10] on NeRF Synthetic dataset in terms of SSIM, PSNR, LPIPS. “Ours” is with NeuS mesh and “Ours(S*)” is with the same mesh as “SuGaR”. The bests are marked in bold, second bests are underlined. ($\uparrow$ means higher is better, $\downarrow$ means lower is better)
[…]
N.E. SuGaR Ours(S*) Ours N.E. SuGaR Ours(S*) Ours N.E. SuGaR Ours(S*) Ours
Chair 28.15 34.23 35.68 35.38 0.943 0.984 0.987 0.986 0.061 0.013 0.010 0.011
Drums 21.14 25.78 26.15 26.19 0.884 0.948 0.953 0.953 0.120 0.044 0.038 0.039
Ficus 23.82 32.09 35.29 35.40 0.909 0.973 0.986 0.986 0.101 0.025 0.013 0.013
Hotdog 32.67 36.53 37.67 37.49 0.969 0.983 0.986 0.984 0.048 0.022 0.017 0.019
Lego 29.16 35.05 36.41 36.33 0.944 0.979 0.983 0.982 0.074 0.018 0.014 0.015
Material 29.48 28.65 29.51 29.91 0.944 0.941 0.951 0.956 0.063 0.055 0.042 0.046
Mic 29.60 35.39 36.99 37.46 0.952 0.990 0.992 0.992 0.046 0.011 0.008 0.007
Ship 25.01 28.55 30.48 31.01 0.083 0.869 0.889 0.890 0.194 0.124 0.097 0.097
Average 32.06 33.52 33.65 0.958 0.966 0.966 0.039 0.030 0.031
This binding strategy allows 3DGS to move outside the triangle while preserving the relative position and rotation of the Gaussians after mesh manipulation.
However, following mesh manipulation, not only does the triangle center change but also the triangle shape. With the altered triangle shape, the local Gaussian position and scaling should adjust accordingly. When the triangle enlarges, it is intuitive that the local scaling and position should expand as well: